A motorcycle starts from rest and accelerates along a straight path at 2 metre per second square. At the starting location of the motorcycle there is a station electric siren. how far has the motorcycle gone when the driver hears the frequency of the siren at 94% of its value when the motorcycle was addressed the speed of the sound is 330 metre per second?1)49m2)147m3)98m4)196m

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Doppler's effect of frequency of source of sound when source is stationary and the observer is moving away from source of sound is given by:
ν' = ν ×(v-vo)/v
 
We are given ν' =90/100 ν, v = 300m/s and v0 = 2√s
Therefore 90/100 ν = ν x(300−2√s)/300
or 270 = 300−2√s
or 2√s = 30
or s = 225m